Grade 12th passPhysical Chemistry

The atomic radius of CR is .1249nm and it has a bcc structure . what is the free volume per unit cell?

komalpreet
10 Years agoGrade 12th pass

1 Answer

Sakshi
10 Years ago
hey student..
for this question you need to use the basic general formula for the body centred cubic unit cell
The relationship between a and R can be worked out by the Pythagorean theorem:(4 R)2 = 3 a2
Thus,
4 R = sqrt(3) a
or
a = 4R/sqrt(3)
Well, this means you should calculate the percentage of the volume occupied by the spheres in the packing. Packing fraction = Vsphere / Vunit cell
= 2*4/3 p R3 / (4/sqrt(3) R3)
= Ö3 p / 8 = 0.6802
nou u know the packing fraction, subtracting it from 1 you can get the empty space also